# DeCoF: Generated Video Detection via Frame Consistency

### v2

@@ 5.2 Shortcomings of learning from spatial artifacts
Figure 8: Overview of the DeCoF framework. We first get real video and generated video features by the pre-trained model CLIP:VIT, which is designed to eliminate the impact of spatial artifacts on capture temporal artifacts, and then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of real and fake video.
5.3 Eliminate impact of spatial artifacts
There is no doubt that real video frames are continuous, and we can even predict any frame of the video through other frames. However, mismatches are prone to occur between the generated video frames, resulting in the inability to meet the continuity of the video frames, that is, temporal artifacts of the generated video. Given a set of video frames $V=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a particular function that measures the continuity between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video.
After analyzing the results of the above exploration, it appears that the detection model tends to learn spatial artifacts that are easy to capture but do not have generalization properties. The difficulty lies in how to eliminate the impact of spatial artifacts on capturing temporal artifacts.
@@ 5.3 Eliminate impact of spatial artifacts
Intuitively, if we can find a particular mapping function $g$ with the following properties: The distance between frame mappings is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts. In fact, we choose ViT-L/14  (Dosovitskiy et al., 2020) trained for the task of image-language alignment as mapping function $g$, then by mapping the video frames $F_{i}$ to $g(F_{i})$, the decoupling of temporal artifacts and spatial artifacts can be achieved. As a result, the original problem can be expressed as $f(g(F_{1}),g(F_{2}),\cdots,g(F_{L}))=y$.
@@
Model framework. In order to learn video frame consistency for capturing temporal artifacts, we design our detection with the Consistency of video Frames (DeCoF), in which the critical verification module consists of two transformer layers and an MLP head, As shown in Figure  8. More specifically, given a video input $V$, The frozen parameter image encoder (ViT-L/14) will extract the features of the video frames and generate a feature sequence $S=\{g(F_{1}),g(F_{2}),\cdots,g(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature. Next, the position encoding and a learnable class embedding representing the classification output are prepended to the sequence $S$ and then passed through a sequential module composed of two layers of transformers. Finally, a classification header is appended to the transformer layer, resulting in a probability distribution over the target class, and we train our model via a simple cross-entropy loss.
@@
In addition to decoupling temporal and spatial artifacts, our approach significantly reduces computational complexity and memory requirements since we only need to learn anomalies between video frame features. Overall, our approach is simple yet effective.
@@ 6 Experiments
In this section, we introduce the experiment setup and provide extensive experimental results to demonstrate the superiority of our approach.
Table 2: Comprehensive comparison of our method with previous detectors and spatiotemporal convolutional networks, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model. * Denotes our retraining on a subset of generated video forensics with official code. We report ACC (%) and AP (%) (ACC and AP in the Table).
Method Training dataset Generate mode Detection Level Tested generate video models Total Avg.
@@
CNNDet(0.5) LSUN ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
CNNDet(0.1) LSUN ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
DIF SDV2 Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
DIRE LSUN-B Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
Lgard LSUN ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
@@
6.1 Experimental Setup
Data pre-processing.We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalization of the detector. We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
Evaluation metrics. Following previous methods  (Wang et al., 2020, 2023b; Frank et al., 2020), we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors. The threshold for computing accuracy is set to 0.5 following  (Wang et al., 2020, 2023b). For image-based methods, we average frame-level predictions into corresponding video-level predictions.
Baseline.  (i) CNNDet (Wang et al., 2020) proposes a model for detecting images generated by CNN. Through a large amount of data enhancement, the model has generalization on various images generated by CNN.  (ii) DIF  (Sinitsa & Fried, 2023) proposes a few-shot learning method for image synthesis.  (iii) DIRE (Wang et al., 2023b) proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models.  (iv) Lgard (Tan et al., 2023) used a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts.  (v) I3D (Carreira & Zisserman, 2017) introduce a new Inflated 3D ConvNet that is based on 2D ConvNet inflation.  (vi) Slow (Feichtenhofer et al., 2019) proposed a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames.  (vii) X3D (Feichtenhofer, 2020) designed a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy.  (viii) Mvit  (Fan et al., 2021) combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition.
6.2 Detailed Evaluation
We conduct comprehensive and challenging experiments to answer the following questions:
Figure 9: Robustness to unseen perturbations. The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression. We test the detectors trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
@@ 6.2 Detailed Evaluation
Q1. How does our detector perform at detecting unseen T2V models compared to previous detectors?
@@
As mentioned in Section 5.2, we found that previous detectors exhibit extraordinary capabilities when detecting specific generative models but lack generalization when unseen generative models are present. In the table 2, we quantitatively evaluate CNNDet (Wang et al., 2020), DIF (Sinitsa & Fried, 2023), DIRE (Wang et al., 2023b), Lgard (Tan et al., 2023), using the pre-trained weights provided by the official repository. Although these detectors have surprising performance in detecting generated images, their performance drops significantly on the task of detecting generated videos, with both ACC and AP below 60%. As a result of retraining the above four detectors on the sub-dataset Text2Video-zero, their performance on specific video generation models greatly improved, but they still performed poorly for unseen generation models, and ACC and AP scores are still lower than 60%, and even the retrained spatiotemporal convolutional network displays similar results. On the contrary, compared with previous detectors, our generated video detection method has excellent generalization for various unseen generation models. Specifically, on the task of generating video detection, the ACC increased by 17.74% $\sim$ 31.38% and the AP by 16.73% $\sim$ 49.18%.
@@
We also conducted cross-dataset evaluations. Table  2 shows that on the task of generating video detection, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%. This is because Text2Video-zero has the worst temporal continuity in all four T2V generation models.
@@
Training dataset Tested generate video models Total Avg.
Gen-2 Pika
ACC AP ACC AP ACC AP
Text2Video-zero 99.49 100.00 99.49 99.98 99.49 99.99
Show1 98.46 99.96 98.46 99.85 98.46 99.91
ModelScope 98.97 99.99 96.89 99.73 97.93 99.86
ZeroScope 98.97 99.99 94.84 99.52 96.91 99.76
In practical detection scenarios, the robustness of the detector to unseen perturbations is also crucial. Here, we mainly focus on the impact of two perturbations on the detector: Gaussian blur and JPEG compression. The perturbations are added under three levels for Gaussian blur ($\sigma$ = 1, 2, 3) and five levels for JPEG compression (quality = 90, 80, 70, 60, 50). Figure  9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations. Although the model’s performance in the face of perturbations is inconsistent due to differences in training data sets, Our method remains effective under any degree of arbitrary perturbation, and the AP of models trained on any sub-dataset is greater than 80%.
Q3. How does our detector perform on commercial generation models?
To further evaluate the effectiveness of our detection method, we constructed two additional test datasets using two well-known commercial proprietary T2V generation models, Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023). Based on the datasets generated by the four open-source T2V models shown in Table  3, the detector was trained on Text2Video-zero, Show1, ModelScope, and ZeroScope. When confronted with the commercial generation model, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2, demonstrating the strong generalization capability of our model DeCoF for unseen T2V models. Even when trained on specific T2V models, the model is highly accurate in detecting newly generated video.
@@ 7 Conclusions and Future Work
The paper presents a groundbreaking investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content. We construct the first public dataset for real and fake video detection, providing a solid benchmark for community research. At the same time, we propose a simple yet effective detection model, which has proven its strong generalizability and robustness on multiple open-source and commercial closed-source models. We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
8 Impact Statements
Our research focuses on using Machine Learning to detect generated videos. We developed the DeCoF model and shared our dataset to support detecting synthetic videos. These works are crucial to protect digital content and prevent the spread of disinformation. Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology. We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity. We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
Appendix A $GeneratedVideoForensics$ Dataset
A.1 Dataset Composition
@@
The following section provides a more detailed description of how the generated video forensics dataset was constructed and evaluated. The entire GVF dataset is generated in two steps:
@@
1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  (Chen & Dolan, 2011) and MSR-VTT  (Xu et al., 2016) datasets. When selecting samples, we followed the work of FETV  (Liu et al., 2023), Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios. More specifically, in the work of FETV, they classify each prompt controlling video generation from two aspects: the main content it describes and the attributes it controls during video generation. The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations. The corresponding spatial attribute controls are divided into quantity, color, and camera view. Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
@@
2) Selection of generation models: Although there are many high-quality generative models on the market, we do not want to choose the best. On the contrary, we have selected four generative models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of Generative models appear, how do we deal with the problems caused by invisible generative models? In the end, the four most representative generative models were selected as our generative models: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a),ZeroScope(Sterling., June 2023),Show1 (Zhang et al., 2023). At the same time, we also selected the two most popular models as the secret test dataset: Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023).
@@
If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$. If divided by the generation model, the GenerateVideoForensics data set can be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and SHOW-1. Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
A.2 Dataset Evaluation

### v3

@@ 5.2 Capturing Universal Spatial Artifacts is Difficult
As shown in Figure 6, on the one hand, the detector has successfully classified the real video and the fake video generated by the seen generate model (Text2Video-zero), indicating that the detector has fully captured the spatial artifacts of the videos generated by the seen generate model. On the other hand, the detector cannot distinguish fake videos generated by unseen models from the real videos. Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting videos generated by other video generation models? We illustrate this by visualizing the frequency spectrum of video frames generated by the models. As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent. Therefore, capturing universal spatial artifacts appears to be extremely difficult.
@@
Figure 8: Overview of the DeCoF framework. We first get real video and generated video features using the pre-trained CLIP:VIT, to eliminate the impact of spatial artifacts on capturing temporal artifacts. Then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of the real and fake videos.
5.3 Detection based on Temporal Artifacts
The recent video generation models convert the T2I model into the T2V model through additional constraints, which have common drawbacks in controlling video continuity in two aspects: 1) the normativity of motion, not only do visible movements violate to physical laws, but more importantly, there are irregular movements of invisible detail pixels, 2) there is a slight pixel shift in the relatively static local background. Therefore, we propose to develop a detection method based on temporal artifacts.
Given a set of video frames $F=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a specific function that measures the consistency between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video. It appears that the detection model tends to learn spatial artifacts (Section 5.1). Therefore, we propose to eliminate the impact of spatial artifacts in order to help the model focus on capturing temporal artifacts.
@@ 5.3 Detection based on Temporal Artifacts
Intuitively, we want to find a particular mapping function $M$ with the following properties: the distance between mapping frames is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts. Hence, we choose ViT-L/14 [8] trained for the task of image-language alignment as mapping function $M$. Then by mapping the video frames $F_{i}$ to $M(F_{i})$, We have implemented decoupling the temporal and spatial artifacts because the mapping function is not trained to capture spatial artifacts, but only encodes the image into a feature space with excellent properties, where the distance between features is inversely proportional to the image similarity. As a result, the original problem can be expressed as $f(M(F_{1}),M(F_{2}),\cdots,M(F_{L}))=y$.
@@
Model framework. In order to learn video frame consistency for capturing temporal artifacts, we design a detection method based on the Consistency of video Frames (DeCoF), in which the core verification module consists of two transformer layers and an MLP head as shown in Figure 8. More specifically, given a video input $V$, the frozen parameter image encoder (ViT-L/14) extracts the features of the video frames and generates a feature sequence $S=\{M(F_{1}),M(F_{2}),\cdots,M(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature. Next, the position encoding and a learnable class embedding representing the classification output are prepended to the sequence $S$ and then passed through a sequential module composed of two layers of transformers. Finally, a classification header is appended to the transformer layer, outputting a probability distribution over the target class. We train our model via a simple cross-entropy loss.
@@
In addition to decoupling temporal and spatial artifacts, our approach significantly reduces computational complexity and memory requirements since we only need to learn anomalies between video frame features. As such, we achieve a simple yet effective approach.
@@ 6 Experiments
6.1 Experimental Setup
Data pre-processing. We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalizability of the detector. We take first 8 frames of each video and center-crop each frame according to its shortest side, then resize it to 224 × 224. During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur are also used.
Evaluation metrics. Following previous methods [47, 48, 13], we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors. The threshold for computing accuracy is set to 0.5 following [47, 48]. For image-based methods, we average frame-level predictions into corresponding video-level predictions.
Baseline. (i) CNNDet [47] detects images generated by GAN. Through a large amount of data enhancement, the model has generalizability on images generated by various GANs. (ii) DIF [39] is a few-shot learning method for image synthesis. (iii) DIRE [48] proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models. (iv) Lgard [43] uses a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts. (v) I3D [5] introduces a new Inflated 3D ConvNet that is based on 2D ConvNet inflation. (vi) Slow [12] proposes a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames. (vii) X3D [11] designs a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy. (viii) Mvit [10] combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition. (ix) UFDetect(UniversalFakeDetect) [29] detects generated images through pretrained CLIP:ViT visual encoders and nearest neighbors. (x) EVR [24] uses the CLIP backbone with frozen parameters and multi layer spatiotemporal Transformer decoder for video action recognition.
6.2 Evaluation
We conduct comprehensive evaluations to answer the following questions:
Table 2: Comprehensive comparison of our method with previous detectors and STCNN, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model. * Denotes we retrained it on a subdataset of GVF dataset with official code. We report ACC (%) and AP (%) (ACC and AP in the Table).
Method Training dataset Generation mode Detection Level Video generation models Total Avg.
@@ 6.2 Evaluation
CNNDet(0.5) - ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
CNNDet(0.1) - ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
DIF - Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
DIRE - Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
Lgard - ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
UFDetect - ProGAN Image 50.13 63.99 53.09 74.42 52.13 75.55 58.89 84.27 53.56 74.56
@@
Q1. How does our detector generalize to unseen T2V models compared to previous detectors?
@@
Table 2 compares CNNDet [47], DIF [39], DIRE [48], Lgard [43], UFDetect [29], using the pre-trained weights provided by the official repository. Although these detectors have good performance in detecting generated images, their performance drops significantly on the task of detect generated videos, with both ACC and AP below 60%. As a result of retraining the above five detectors on the sub-dataset Text2Video-zero, their performance on seen video generation models greatly improved, but they still performed poorly for unseen generation models, with ACC and AP scores still lower than 60%, and the retrained STCNN presents similar results. On the contrary, our DeCoF has excellent generalizability for various unseen generation models. Compared with previous video classification methods based on pretrained CLIP [24], our method has still made significant progress. Specifically, on the task of generated video detection, DeCoF outperforms previous detectors by 13.77% $\sim$ 31.38% in ACC and 8.44% $\sim$ 49.18% in AP.
@@
We also conducted cross-dataset evaluations. As shown in Table 2, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%.
Figure 9: Robustness to unseen perturbations. The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression. We test the detector trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
In practical detection scenarios, the robustness of the detector to unseen perturbations is also crucial. Here, we mainly focus on the impact of two perturbations on the detector: Gaussian blur and JPEG compression. The perturbations are added under three levels for Gaussian blur ($\sigma$ = 1, 2, 3) and five levels for JPEG compression (quality = 90, 80, 70, 60, 50). Figure 9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations. Although experiencing some performance inconsistencies due to differences between training datasets, our method remains effective under all degrees of arbitrary perturbation, with the AP higher than 80%.
Q3. How does our detector perform on commercial generation models?
@@
Training dataset Video generation models
Gen-2 Pika Sora Veo
ACC AP ACC AP ACC AP ACC AP
Text2Video-zero 99.49 100.00 99.49 99.98 96.87 99.70 96.43 100.00
Show1 98.46 99.96 98.46 99.85 94.79 98.26 92.86 100.00
ModelScope 98.97 99.99 96.89 99.73 94.79 98.83 89.29 98.47
ZeroScope 98.97 99.99 94.84 99.52 89.58 97.92 67.86 94.90
To further evaluate the effectiveness of our detection method, we constructed four additional test datasets using well-known commercial proprietary T2V models, Pika [32], Gen-2 [15], Sora [41] and Veo [45]. As shown in Table 3, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2. Due to regional restrictions, we did not qualify for testing Sora and Veo directly, but we constructed test sets for 48 public videos released by Sora and 14 public videos released by Veo, respectively. In this test, we achieved no less than 89.58% ACC and no less than 97.92% AP on Sora, no less than 67.86% ACC and no less than 94.90% AP on Veo, demonstrating the strong generalizability of our model DeCoF for unseen commercial T2V models.
@@ 7 Conclusions and Future Work
The paper presents the first investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content. We construct the first public dataset for video detection, providing a solid benchmark for community research. At the same time, we propose a simple yet effective detection model, and prove its strong generalizability and robustness on multiple open-source and commercial closed-source models. We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
Appendix A Appendix
This supplementary material provides:
  • Sec. B: A detailed description of dataset construction, including dataset structure diagrams, as well as specific examples of prompts and videos.
  • Sec. C : Various parameter details for model training, hyperparameter settings, etc.
  • Sec. D : The impact of selecting different mapping functions.
  • Sec. E : The impact of video frame count.
  • Sec. F : The impact of video frame sampling interval.
  • Sec. G : More details on Sora.
  • Sec. H : Spatial artifacts vs Temporal artifacts.
  • Sec. I : Impact Statements.
  • Sec. J : Limitation.
  • Sec. K : More visual examples.
Appendix B $GeneratedVideoForensics$ Dataset
B.1 Dataset Composition
@@
The following section provides a more detailed description of how the $GeneratedVideoForensics$ was constructed and evaluated. The entire GVF dataset is generated in two steps:
@@
1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  [6] and MSR-VTT  [50] datasets. When selecting samples, we followed the work of FETV  [25], Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios. More specifically, in the work of FETV, they classify each prompt from two aspects: the main content it describes and the attributes it controls during video generation. The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations. The corresponding spatial attribute controls are divided into quantity, color, and camera view. Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
@@
2) Selection of generation models: Although there are many high-quality generation models on the market, we do not want to choose the best. On the contrary, we have selected four generation models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of generation models appear, how do we deal with the problems caused by unseen generation models? In the end, the four most representative generation models were selected as our generation models: Text2Video-zero [22], ModelScopeT2V [46], ZeroScope[42], Show1 [51]. At the same time, we also selected the three most popular models as the secret test dataset: Pika [32], Gen-2 [15], Sora[41] and Veo [45]. For Sora and Veo, we provide detailed information in the section G
@@
If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$.
If divided by the generation model, the GenerateVideoForensics data set can be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1. Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
B.2 Dataset Evaluation